OpenStudy (anonymous):

approximately how many times difference in the concentration of H+ is there between water and orange juice?

OpenStudy (aaronq):

is there any data to go with that?

OpenStudy (anonymous):

i'm going to guess that orange juice is a buffer of citrate acid and we are around the second half-equivalence point, so the pH is about the same as the pKa2, which is about 4.76 "how many times different are thing 1 and thing 2" can be translated into "what is the ratio of thing 1 to thing 2", so we want the ratio of H plus concentrations. therefore: Hplus (water) / Hplus (orange) = 10^(-pH of water) / 10^(-pH of orange) = 10^(-7) / 10^(-4.76) = 0.005 So the H plus concentration in water is about half of a percent of the H plus concentration in orange.

OpenStudy (anonymous):

it makes sense because we can guess that orange juice is acidic (because it tastes sour, according to an archaic definition of acidity). i made many guesses here, including the fact that the temperature is around 25 deg Celsius (or the pH of water won't be 7). in any case, it sounds like they just want an estimate, so it's ok to guess.